# %TODO␣FINAL:␣Replace␣with␣your␣institution␣list.www.retrocausal.ai

### v6

@@ 0.A.1 Implementation Details
Hyperparameter Settings. Tab. 10 includes a summary of our hyperparameter settings. We follow the same hyperparameter settings used in CASA [43], e.g., frames per second, batch size, temperature, learning rate, and weight decay. We keep the augmentation probability the same as CASA [43]. However, we adjust the noise standard deviation to 0.1 radian for orientation augmentation and 1 heatmap pixel for position augmentation, since our augmentation techniques are conducted on 2D skeletons/heatmaps. For CASA [43], the input vector length increases linearly with the number of keypoints and the number of objects (i.e., humans, hands) involved in each dataset. In contrast, our approach relies on the same heatmap spatial dimension of 30 $\times$ 30 for all datasets, demonstrating better scalability. In addition, for our self-attention module, we follow the same hyperparameter settings used in ViViT [2], except that we reduce the patch size to 10 $\times$ 10 and the transformer output dimension to 192 since our heatmap spatial dimension is small (i.e., 30 $\times$ 30).
@@ 0.A.3 Multi-Modality Fusion
We first illustrate our multi-modality model architecture in Fig. 6, which we describe in details in Sec. 3.4 of the main paper. Next, our complete multi-modality fusion results on all metrics and datasets are presented in Tab. 12. It is clear from the results, our RGB+Pose model yields the best numbers across all metrics and datasets, outperforming all the single-modality methods and the multi-modality baseline extended from CASA [43] and establishing the state-of-the-art across all metrics and datasets. In addition, our Pose-only model obtains better results than our RGB-only model on Penn Action and H2O, while having generally worse results on IKEA ASM. It likely implies that pose cues are more important on Penn Action and H2O, while RGB cues are more crucial on IKEA ASM. Finally, both our RGB+Pose model and the CASA-based multi-modality baseline outperform their single-modality counterparts respectively, which validates the benefit of multi-modality inputs for fine-grained human activity understanding tasks.
@@
Penn Action LAV [30] $\square$ 83.56 83.95 84.25 0.6613 0.8047 79.13 78.98 78.90
VAVA [50] $\square$ 83.89 84.23 84.48 0.7091 0.8053 81.52 80.47 80.67
CASA [43] $\blacktriangle$ 88.55 91.87 92.20 0.9449 0.9728 89.90 89.44 89.07
CASA [43] $\blacktriangle$ + $\square$ 88.97 92.55 92.90 0.9494 0.9808 91.74 90.97 90.25
@@
IKEA ASM LAV [30] $\square$ 29.78 29.85 30.43 - - 23.89 23.65 23.56
VAVA [50] $\square$ 31.66 33.79 32.91 - - 29.58 28.74 28.48
CASA [43] $\blacktriangle$ 21.32 31.52 31.06 - - 28.92 28.88 28.61
CASA [43] $\blacktriangle$ + $\square$ 27.63 32.85 34.33 31.45 31.08 30.87
@@
H2O LAV [30] $\square$ 35.38 51.66 53.43 0.5913 0.5323 47.55 45.56 44.61
CASA [43] $\blacktriangle$ 43.50 62.51 68.78 0.9107 0.9438 60.13 59.44 59.01
@@ 0.A.4 Effects of Different Method Parts
To analyze sources of performance gains, we conduct an ablation study where various parts of our method, i.e., 2D skeleton heatmap inputs + 2D skeleton heatmap augmentations, spatial transfer, and multi-modality fusion, are enabled. Tab. 13 presents the results of CASA [43] (top row) and our variants. In particular, the second row shows the results of our variant ($\heartsuit$), where we disable our 2D skeleton heatmap inputs+2D skeleton heatmap augmentations (and use the 3D skeleton coordinate inputs+3D skeleton coordinate augmentations of CASA [43] instead) and our multi-modality fusion, and further replace the spatial image transformer of ViViT [2] with the spatial graph transformer of ST-TR [60] due to 3D skeleton coordinate inputs. Next, the third row shows the results of our variant ($\diamondsuit$), where we enable our 2D skeleton heatmap inputs+2D skeleton heatmap augmentations, but disable our spatial transformer (and use MLPs instead) and our multi-modality fusion. Lastly, the forth row shows the results of our variant without multi-modality fusion ($\spadesuit$), while the bottom row shows the results of our complete method ($\clubsuit$).
From the results in Tab. 13, using either 2D skeleton heatmap inputs+2D skeleton heatmap augmentations ($\heartsuit$) or spatial transformer ($\diamondsuit$) yields minor performance gains over CASA [43]. Moreover, using both 2D skeleton heatmap inputs+2D skeleton heatmap augmentations and spatial transformer ($\spadesuit$) leads to major performance gains over CASA [43]. Finally, adding multi-modality fusion ($\clubsuit$) boosts our results further.
@@
H2O CASA [43] $\times$ $\times$ $\times$ 68.78 0.9107 0.9438
@@ 0.A.5 Effects of Different 2D Poses
In all of the previous experiments, we have used 2D poses that are projected (via orthographic projection) from 3D poses provided by CASA [43] for fair comparison purposes. We use orthographic projection since camera intrinsics/extrinsics are not available for all datasets (e.g., Penn Action). In this section, we study the impacts of different 2D poses on the performance of our approach. Specifically, we use estimated 2D poses (computed by OpenPose [8]) on Penn Action and IKEA ASM and ground truth 2D poses on IKEA ASM (ground truth 2D poses are not available on Penn Action). The ablation results of different 2D poses are shown in Tab. 14. It can be seen that that using OpenPose 2D poses generally yields better performance than using our projected 2D poses and further advances the state-of-the-art in this topic. This is likely because our projected 2D poses have deformations/deviations due to orthographic projection. Examples of our projected 2D poses are plotted in Fig. 7. Furthermore, using ground truth 2D poses yields the best results, outperforming using OpenPose 2D poses and using our projected 2D poses.
@@ 0.A.6 Qualitative Results
Firstly, Figs. 8 and 9 illustrate the alignment results of our approach and CASA [43] between two bowling sequences and two pullups sequences. It is evident from the figures that LA2DS achieves more seamless alignment results than CASA [43]. Our method learns interactions between joints in the same skeleton and between skeletons in the same sequence and across the sequences to align them in time effectively.
Moreover, Fig. 10 shows the fine-grained frame retrieval results of our approach, CASA [43], and LAV [30] on clean_and_jerk and jumping_jacks sequences. LA2DS extracts useful spatiotemporal and contextual features within each sequence, resulting in more accurate fine-grained frame retrieval results than CASA [43] and LAV [30], as can be seen in Fig. 10.
Finally, we show the alignment results, i.e., framewise correspondences, between two tennis_serve sequences in Fig. 11, and the t-SNE [53] visualization of the learned embeddings of two baseball_pitch sequences in Fig. 12. Our method achieves a more seamless alignment as compared to CASA [43].
@@ 0.A.8 Limitations
Challenging Viewpoints and Extreme Viewpoint Gaps. We follow all of previous works to use Penn Action, IKEA ASM, and H2O, which contain frontal, side, or top viewpoints of human actions (e.g., Fig. 5 of the main paper), since it is difficult to capture human actions with rear or bottom viewpoints due to severe occlusions or privacy concerns respectively. Also, it is generally harder to obtain accurate 3D poses than accurate 2D poses, especially for challenging viewpoints, thus CASA [43]’s sensitivity to noise/errors of pose inputs may further be amplified with challenging viewpoints. However, we do randomly pair videos to perform alignment, yielding viewpoint gaps of up to $180$ degrees (e.g., pairing videos with left and right side viewpoints in Figs. 5(a-b) of the main paper). Collecting a human action dataset including rear or bottom viewpoints and learning by aligning videos with extreme viewpoint gaps are challenging and remain our future works.
Missing Context Details. Both 3D skeleton inputs (e.g., CASA [43]) and 2D skeleton inputs (i.e., our method) do not include context details, e.g., objects and background information, which are available in RGB video inputs. Thus, our method may fail in cases where context details are important in recognizing the actions. Nevertheless, our multi-modality fusion method can overcome this limitation by utilizing both 2D skeleton heatmaps and RGB videos as inputs.

### v7

@@ 0.A.1 Implementation Details
Hyperparameter Settings. Tab. 10 includes a summary of our hyperparameter settings. We follow the same hyperparameter settings used in CASA [kwon2022context], e.g., frames per second, batch size, temperature, learning rate, and weight decay. We keep the augmentation probability the same as CASA [kwon2022context]. However, we adjust the noise standard deviation to 0.1 radian for orientation augmentation and 1 heatmap pixel for position augmentation, since our augmentation techniques are conducted on 2D skeletons/heatmaps. For CASA [kwon2022context], the input vector length increases linearly with the number of keypoints and the number of objects (i.e., humans, hands) involved in each dataset. In contrast, our approach relies on the same heatmap spatial dimension of 30 $\times$ 30 for all datasets, demonstrating better scalability. In addition, for our self-attention module, we follow the same hyperparameter settings used in ViViT [arnab2021vivit], except that we reduce the patch size to 10 $\times$ 10 and the transformer output dimension to 192 since our heatmap spatial dimension is small (i.e., 30 $\times$ 30).
@@ 0.A.3 Multi-Modality Fusion
We first illustrate our multi-modality model architecture in Fig. 6, which we describe in details in Sec. 3.4 of the main paper. Next, our complete multi-modality fusion results on all metrics and datasets are presented in Tab. 12. It is clear from the results, our RGB+Pose model yields the best numbers across all metrics and datasets, outperforming all the single-modality methods and the multi-modality baseline extended from CASA [kwon2022context] and establishing the state-of-the-art across all metrics and datasets. In addition, our Pose-only model obtains better results than our RGB-only model on Penn Action and H2O, while having generally worse results on IKEA ASM. It likely implies that pose cues are more important on Penn Action and H2O, while RGB cues are more crucial on IKEA ASM. Finally, both our RGB+Pose model and the CASA-based multi-modality baseline outperform their single-modality counterparts respectively, which validates the benefit of multi-modality inputs for fine-grained human activity understanding tasks.
@@
Penn Action LAV [haresh2021learning] $\square$ 83.56 83.95 84.25 0.6613 0.8047 79.13 78.98 78.90
VAVA [liu2022learning] $\square$ 83.89 84.23 84.48 0.7091 0.8053 81.52 80.47 80.67
CASA [kwon2022context] $\blacktriangle$ 88.55 91.87 92.20 0.9449 0.9728 89.90 89.44 89.07
CASA [kwon2022context] $\blacktriangle$ + $\square$ 88.97 92.55 92.90 0.9494 0.9808 91.74 90.97 90.25
@@
IKEA ASM LAV [haresh2021learning] $\square$ 29.78 29.85 30.43 - - 23.89 23.65 23.56
VAVA [liu2022learning] $\square$ 31.66 33.79 32.91 - - 29.58 28.74 28.48
CASA [kwon2022context] $\blacktriangle$ 21.32 31.52 31.06 - - 28.92 28.88 28.61
CASA [kwon2022context] $\blacktriangle$ + $\square$ 27.63 32.85 34.33 31.45 31.08 30.87
@@
H2O LAV [haresh2021learning] $\square$ 35.38 51.66 53.43 0.5913 0.5323 47.55 45.56 44.61
CASA [kwon2022context] $\blacktriangle$ 43.50 62.51 68.78 0.9107 0.9438 60.13 59.44 59.01
@@ 0.A.4 Effects of Different Method Parts
To analyze sources of performance gains, we conduct an ablation study where various parts of our method, i.e., 2D skeleton heatmap inputs + 2D skeleton heatmap augmentations, spatial transfer, and multi-modality fusion, are enabled. Tab. 13 presents the results of CASA [kwon2022context] (top row) and our variants. In particular, the second row shows the results of our variant ($\heartsuit$), where we disable our 2D skeleton heatmap inputs+2D skeleton heatmap augmentations (and use the 3D skeleton coordinate inputs+3D skeleton coordinate augmentations of CASA [kwon2022context] instead) and our multi-modality fusion, and further replace the spatial image transformer of ViViT [arnab2021vivit] with the spatial graph transformer of ST-TR [plizzari2021skeleton] due to 3D skeleton coordinate inputs. Next, the third row shows the results of our variant ($\diamondsuit$), where we enable our 2D skeleton heatmap inputs+2D skeleton heatmap augmentations, but disable our spatial transformer (and use MLPs instead) and our multi-modality fusion. Lastly, the forth row shows the results of our variant without multi-modality fusion ($\spadesuit$), while the bottom row shows the results of our complete method ($\clubsuit$).
From the results in Tab. 13, using either 2D skeleton heatmap inputs+2D skeleton heatmap augmentations ($\heartsuit$) or spatial transformer ($\diamondsuit$) yields minor performance gains over CASA [kwon2022context]. Moreover, using both 2D skeleton heatmap inputs+2D skeleton heatmap augmentations and spatial transformer ($\spadesuit$) leads to major performance gains over CASA [kwon2022context]. Finally, adding multi-modality fusion ($\clubsuit$) boosts our results further.
@@
H2O CASA [kwon2022context] $\times$ $\times$ $\times$ 68.78 0.9107 0.9438
@@ 0.A.5 Effects of Different 2D Poses
In all of the previous experiments, we have used 2D poses that are projected (via orthographic projection) from 3D poses provided by CASA [kwon2022context] for fair comparison purposes. We use orthographic projection since camera intrinsics/extrinsics are not available for all datasets (e.g., Penn Action). In this section, we study the impacts of different 2D poses on the performance of our approach. Specifically, we use estimated 2D poses (computed by OpenPose [cao2019openpose]) on Penn Action and IKEA ASM and ground truth 2D poses on IKEA ASM (ground truth 2D poses are not available on Penn Action). The ablation results of different 2D poses are shown in Tab. 14. It can be seen that that using OpenPose 2D poses generally yields better performance than using our projected 2D poses and further advances the state-of-the-art in this topic. This is likely because our projected 2D poses have deformations/deviations due to orthographic projection. Examples of our projected 2D poses are plotted in Fig. 7. Furthermore, using ground truth 2D poses yields the best results, outperforming using OpenPose 2D poses and using our projected 2D poses.
@@ 0.A.6 Qualitative Results
Firstly, Figs. 8 and 9 illustrate the alignment results of our approach and CASA [kwon2022context] between two bowling sequences and two pullups sequences. It is evident from the figures that LA2DS achieves more seamless alignment results than CASA [kwon2022context]. Our method learns interactions between joints in the same skeleton and between skeletons in the same sequence and across the sequences to align them in time effectively.
Moreover, Fig. 10 shows the fine-grained frame retrieval results of our approach, CASA [kwon2022context], and LAV [haresh2021learning] on clean_and_jerk and jumping_jacks sequences. LA2DS extracts useful spatiotemporal and contextual features within each sequence, resulting in more accurate fine-grained frame retrieval results than CASA [kwon2022context] and LAV [haresh2021learning], as can be seen in Fig. 10.
Finally, we show the alignment results, i.e., framewise correspondences, between two tennis_serve sequences in Fig. 11, and the t-SNE [tsne2008] visualization of the learned embeddings of two baseball_pitch sequences in Fig. 12. Our method achieves a more seamless alignment as compared to CASA [kwon2022context].
@@ 0.A.8 Limitations
Challenging Viewpoints and Extreme Viewpoint Gaps. We follow all of previous works to use Penn Action, IKEA ASM, and H2O, which contain frontal, side, or top viewpoints of human actions (e.g., Fig. 5 of the main paper), since it is difficult to capture human actions with rear or bottom viewpoints due to severe occlusions or privacy concerns respectively. Also, it is generally harder to obtain accurate 3D poses than accurate 2D poses, especially for challenging viewpoints, thus CASA [kwon2022context]’s sensitivity to noise/errors of pose inputs may further be amplified with challenging viewpoints. However, we do randomly pair videos to perform alignment, yielding viewpoint gaps of up to $180$ degrees (e.g., pairing videos with left and right side viewpoints in Figs. 5(a-b) of the main paper). Collecting a human action dataset including rear or bottom viewpoints and learning by aligning videos with extreme viewpoint gaps are challenging and remain our future works.
Missing Context Details. Both 3D skeleton inputs (e.g., CASA [kwon2022context]) and 2D skeleton inputs (i.e., our method) do not include context details, e.g., objects and background information, which are available in RGB video inputs. Thus, our method may fail in cases where context details are important in recognizing the actions. Nevertheless, our multi-modality fusion method can overcome this limitation by utilizing both 2D skeleton heatmaps and RGB videos as inputs.